OpenStudy (anonymous):

An object of mass 3.8 kg is pushed directly from rest along a horizontal surface, a distance of 1.2, and reaches a speed of 2.6 m/s by the end of the push. The frictional force acting is 6.7 N. Determine the acceleration of the object while being pushed.

OpenStudy (roseflight):

As far as I can tell you have a lot of extraneous data. Here's how I would solve it: d=1.2 (assuming these are meters) Vi=0 m/s Vf=2.6 m/s a=? m/s^2 Use \[Vf^{2}=Vi ^{2}+2ad\] \[(2.60^{2}=(0^2)+2(a)1.2)\]